# Applications of Automatic Differentiation in Image Registration

### v1

@@ Abstract
We demonstrate that automatic differentiation, which has become commonly available in machine learning frameworks, is an efficient way to explore ideas that lead to algorithmic improvement in multi-scale affine image registration and affine super-resolution problems. In our first experiment on multi-scale registration, we implement an ODE predictor-corrector method involving a derivative with respect to the scale parameter and the Hessian of an image registration objective function, both of which would be difficult to compute without AD. Our findings indicate that exact Hessians are necessary for the method to provide any benefits over a traditional multi-scale method; a Gauss-Newton Hessian approximation fails to provide such benefits. In our second experiment, we implement a variable projected Gauss-Newton method for super-resolution and use AD to differentiate through the iteratively computed projection, a method previously unaddressed in the literature. We show that Jacobians obtained without differentiating through the projection are poor approximations to the true Jacobians of the variable projected forward map and explore the performance of some other approximations. By addressing these problems, this work contributes to the application of AD in image registration and sets a precedent for further use of machine learning tools in this field.
@@ 1 Introduction
Leveraging the PyTorch [15] library, we consider two novel applications of automatic differentiation (AD) to this framework. We apply a predictor-corrector method method to perform multi-scale image registration and use variable projection to solve a super-resolution problem involving image reconstruction and the simultaneous registration of multiple images. The prediction step of the predictor-corrector method requires a derivative with respect to a scaling parameter in the interpolation, and a variable projection involves differentiation through an inexact iterative least squares solve. Much of the theory for these problems exists in the literature, but perhaps due to the difficulty of some required derivatives, they do not exist in standard image registration packages like FAIR [12]. The code and data required generate the figures used in this text and to reproduce results can be found at our GitHub repository [19].
2 Background
@@ 2.1 Numerical Optimization Framework
Assume our images are supported on a box-shaped subset $\Omega\subset\mathbb{R}^{d}$ for $d=2$ or $3$ We consider an enumerated set ${\boldsymbol{\omega}}=\{\vec{x}_{1},\dots,\vec{x}_{n}\}\subset\Omega$ of points with even spacing $h_{x}$ and $h_{y}$ in each dimension, on which we have evaluations of the reference and template images. (We do not have evaluations for points not on the grid, which we will have to address.) Using this data, we would like to be able to compute the distance $\mathcal{D}[{\cal T}\circ\vec{y},{\cal R}]$ of our choice, which in this work will be the least-squares distance metric
@@
The final thing to discretize is the transformation $\vec{y}.$ The most expressive “non-parametric” approach would be to solve for every value in $\vec{y}({\boldsymbol{\omega}})$ separately. However, to avoid such a high dimensional optimization problem, we use a parameterization in many fewer parameters ${\bf w}\in\mathbb{R}^{p}$, henceforth writing $\vec{y}(\vec{\cdot};{\bf w})$. A number of standard parameterizations are covered in [12]. In this work, we consider only affine transformations.
@@
(2) $J({\bf w})=\|{\cal T}(\vec{y}({\boldsymbol{\omega}};{\bf w}))-{\cal R}({\boldsymbol{\omega}})\|_{2}^{2}h_{x}h_{y}+\lambda S({\bf w}),$
@@ 2.2 Automatic Differentiation
To help demonstrate the efficiency of AD, Table 1 shows average evaluation times of derivative computations with an affine transformation of $6$ parameters, and a transformation constructed with a Neural ODE [1] [17] [20] of $50$ parameters. The forward function, ${\cal T}(\vec{y}({\boldsymbol{\omega}};{\bf w})$, and loss function $J({\bf w})$ follow (2). Gradient, Hessian, and Jacobian computations are done using Functorch [9] which provides composable function transforms inspired by the Jax Framework [16]. It is important to note that the cost of computing a Hessian with AD grows much quicker than the cost of computing a Jacobian with AD as the number of parameters increases. This suggests that the cost difference between Newton (where Hessians are computed) and Quasi-Newton (where Hessians are approximated) increases as the number of parameters of the transformation increases.
@@
Loss Function $2.171\pm 0.259$ $1$x $5.012\pm 2.168$ $1$x
Gradient $5.422\pm 0.409$ $2.5$x $13.65\pm 1.018$ $2.7$x
Hessian $12.44\pm 1.798$ $5.7$x $120.0\pm 8.737$ $24$x
Forward Function $1.054\pm 0.076$ $1$x $3.204\pm 0.149$ $1$x
Jacobian $11.59\pm 1.518$ $11$x $54.31\pm 3.240$ $17$x
Table 1: Cost of derivative computations using AD. Factor is the approximate constant factor between an evaluation time and the evaluation time of the loss function or forward function (ie. gradient evaluation time took about $2.5$ times the loss function evaluation time for the two examples).
@@ 3.1 Derivation of Multi-scale ODE
Each step of this ODE solve requires an inverse Hessian-vector product, so it is as feasible to implement as Newton’s method.
@@ 3.2 Predictor-Corrector Method
The approach involves taking larger steps in $\theta$ when solving (4), and the error that is produced from that large step can be corrected by performing local minimization. In practice, Newton’s method provides rapid local convergence, and allows us the ability to re-use computed Hessians when taking a step in the ODE solve.
@@
(5) ${\bf w}^{*}_{\rm pred}={\bf w}^{*}(\theta_{n+1})+(\theta_{n}-\theta_{n+1})\left(\frac{d}{d\theta}{\bf w}^{*}(\theta_{n+1})\right)$
of forward Euler’s method is taken to predict the location of ${\bf w}^{*}(\theta_{n+1}),$ and is then corrected by a local minimization procedure to obtain ${\bf w}^{*}(\theta_{n}).$
@@ 3.3 Results
Two results are shown in this section. First, the predictor-corrector method is shown to work for a problem where registering from a single scale does not. This motivates the method as having a similar use case as the traditional multi-scale method. Then, for the predictor-corrector method, the prediction is shown to fail when using approximated Hessians via Gauss Newton, but succeeds when using exact Hessians computed using AD. An implementation of Levenberg-Marquardt is used for Gauss Newton, while Scipy’s Exact Trust Region [18] is used for Newton.
@@ 3.3.1 Single Scale vs Predictor-Corrector
Consider an affine, non-regularized image registration problem with template and reference shown in the first row of Figure 2. As shown in the figure, some image registration problems benefit from using the predictor-corrector method, as opposed to optimizing from a single scale in order to avoid getting stuck in a local minimum.
@@ 3.3.2 Approximated Hessian vs Exact Hessian
To evaluate the effectiveness of the prediction at each scale, i.e., the step of forward Euler 5 , we compare the loss and gradient at the predicted point to the loss and gradient without the prediction. This comparison effectively compares the predictor-corrector method with the traditional multi-scale method. We define the relative loss difference as $(J({\bf w}^{*})-J({\bf w}_{\rm pred}^{*}))/J({\bf w}^{*})$, and the relative gradient norm difference as $(\|\nabla J({\bf w}^{*})\|-\|\nabla J({\bf w}_{\rm pred}^{*})\|)/\|\nabla J({\bf w}^{*})\|$ to give measures of how helpful the predictions are.
This comparison is illustrated in Figure 3. A positive value indicates a successful reduction in loss or gradient, whereas a negative value indicates an increase. The results show that using exact Hessians via Newton’s method tends to decrease the loss and gradient, while using approximated Hessians via Gauss-Newton tends to increase the loss and gradient. Therefore, using exact Hessians is necessary for the predictor-corrector method to provide any benefits over a traditional multi-scale method.
Figure 3: Relative loss and gradient norm, both plotted against $\theta$ on a logarithmic scale. Exact Hessians are used for blue, approximated Hessians are used for orange. All relative gradient norm difference values using approximated Hessians are less than $-31000$.
@@ 4 Coupled Methods for Super-Resolution
In the problem of super-resolution, we aim to reconstruct a high-resolution reference image, $\mathbf{f}^{(0)}$ from a vector of low resolution template images $\mathbf{d}=[{\mathbf{d}^{(0)}}^{\top},\dots,{\mathbf{d}^{(q)}}^{\top}]^{\top}$, which are assumed to be unregistered, down-sampled versions of $\mathbf{f}^{(0)}$. Our aim in super-resolution is to register multiple templates $\mathbf{d}^{(j)}$ to an unknown reference $\mathbf{f}^{(0)},$ while at the same time reconstructing $\mathbf{f}^{(0)}.$
4.1 The Super-Resolution Problem
We follow the framework developed in [3]. Interpolation of the reference is linear in the reference intensity vector ${\bf f}^{(0)}$, thus, we can mathematically model our transformations $\vec{y}^{(j)}$ using interpolation matrices $\mathsf{I}^{(j)}({\bf w}).$ Note that the interpolation matrix is non-linear in its dependence on the transformation parameter vector ${\bf w}.$ Our high resolution transformed templates ${\bf f}^{(j)}\in\mathbb{R}^{n},j=1,\dots,q$ are thus given by the transformations $\mathsf{I}^{(j)}({\bf w}){\bf f}^{(0)}$ of an initial reference image ${\bf f}^{(0)}$. Our observed template images ${\bf d}^{(j)}\in\mathbb{R}^{m},j=0,1\dots,q$ are modeled as ${\bf d}^{(j)}=\mathsf{K}{\bf f}^{(j)},$ where $\mathsf{K}\in\mathbb{R}^{m\times n}$ is a linear operator which lowers the image resolution ($m<n$), in our case by pooling. Note that our modeling assumes that the template ${\bf d}^{(0)}$ is registered with the unknown reference ${\bf f}^{(0)}$, as the data can only ever constrain ${\bf f}^{(0)}$ up to a transformation; registering ${\bf d}^{(0)}$ with ${\bf f}^{(0)}$ eliminates this non-uniqueness and saves some computational overhead. We further let $h^{{\bf f}}_{x},h^{{\bf f}}_{y},h^{{\bf d}}_{x}$ and $h^{{\bf d}}_{y}$ be the corresponding grid sizes for the template and reference grids.
@@ 4.1 The Super-Resolution Problem
$J({\bf w},{\bf f}^{(0)})=\|(\mathsf{I}\otimes\mathsf{K})\mathsf{I}({\bf w}){\bf f}^{(0)}-{\bf d}\|^{2}h^{{\bf d}}_{x}h^{{\bf d}}_{y}+\lambda_{{\bf w}}^{2}S_{{\bf w}}({\bf w})+\lambda_{{\bf f}}^{2}S_{{\bf f}^{(0)}}({\bf f}^{(0)}).$
@@ 4.2 Variable Projection Theory
We begin by noticing that (6) is in the separable form addressed by the variable projection method [7] [6]. To make this even more clear, we write it in the stacked form
$J({\bf w},{\bf f}^{(0)})=\left\|\begin{bmatrix}(\mathsf{I}\otimes\mathsf{K})\mathsf{I}({\bf w})\sqrt{h^{{\bf d}}_{x}h^{{\bf d}}_{y}}\\ \lambda_{{\bf f}}\mathsf{L}\sqrt{h^{{\bf f}}_{x}h^{{\bf f}}_{y}}\end{bmatrix}{\bf f}-\begin{bmatrix}{\bf d}\sqrt{h^{{\bf d}}_{x}h^{{\bf d}}_{y}}\\ \mathsf{O}\end{bmatrix}\right\|^{2}=\|\mathsf{A}_{\lambda_{{\bf f}}}({\bf w}){\bf f}^{(0)}-{\bf b}\|^{2}.$
@@
(7) $\tilde{J}({\bf w})=\|\mathsf{A}_{\lambda_{{\bf f}}}({\bf w}){\bf f}^{(0)}({\bf w})-{\bf b}\|^{2}=\bigg{\|}\rm{Proj}_{\rm{N}(\mathsf{A}_{\lambda_{{\boldsymbol{\alpha}}}}({\bf w})^{\top})}({\bf b})\bigg{\|}^{2}$
@@
Computing the Jacobian $\frac{\partial{\bf f}^{(0)}({\bf w})}{\partial{\bf w}}$ of the projection ${\bf f}^{(0)}({\bf w})$ requires differentiation through the least squares solution. An exact formula for such a derivative based on implicit differentiation through an exact solution exists [7], but it is expensive to compute. Previous work in variable-projected Gauss-Newton has taken the approach of making SVD-based approximations to it [13], or taking advantage of the compute saved with an inexact iterative solution but ignoring $\frac{\partial{\bf f}^{(0)}({\bf w})}{\partial{\bf w}}$ [3].
In an automatic differentiation framework, we can ignore the projection term by turning gradient tracking off during the least squares solve. However, by choosing to track gradients through the least squares solution, we can include this term at the cost of increased memory overhead. It has been shown that automatic differentiation through the Conjugate Gradient (CG) algorithm gives correct gradients in the case of an inexact solve [8], and in the case of an exact solve [2], which we will use in our solutions. A recent preprint [10] also explores automatic differentiation through other least-squares solvers.

### v2

@@ Abstract
We demonstrate that automatic differentiation (AD), which has become commonly available in machine learning frameworks, is an efficient way to explore ideas that lead to algorithmic improvement in multi-scale affine image registration and affine super-resolution problems. In our first experiment on multi-scale registration, we implement an ODE predictor-corrector method involving a derivative with respect to the scale parameter and the Hessian of an image registration objective function, both of which would be difficult to compute without AD. Our findings indicate that exact Hessians are necessary for the method to provide any benefits over a traditional multi-scale method; a Gauss-Newton Hessian approximation fails to provide such benefits. In our second experiment, we implement a variable projected Gauss-Newton method for super-resolution and use AD to differentiate through the iteratively computed projection, a method previously unaddressed in the literature. We show that Jacobians obtained without differentiating through the projection are poor approximations to the true Jacobians of the variable projected forward map and explore the performance of other approximations in the problem of super-resolution. By addressing these problems, this work contributes to the application of AD in image registration and sets a precedent for further use of machine learning tools in this field.
@@ 1 Introduction
The problem of super-resolution, which we properly introduce in §4, seeks to reconstruct an unknown high resolution reference image from a sequence of unregistered low resolution templates [3]. Solution techniques for this problem implement on algorithms for image registration while simultaneously reconstructing the high resolution reference image. We use the same strategy as in registration to solve the super-resolution problem by using a finite dimensional approximation to the problem. However, due to the simultaneous reconstruction of the reference image and the registration of the templates, the objective has a separable structure amenable to a variable projection [9] approach.
Leveraging the PyTorch [18] library, we consider two novel applications of automatic differentiation (AD) to this framework. We apply a predictor-corrector method to perform multi-scale image registration and improve existing variable projection methodology to solve a super-resolution problem. The prediction step of the predictor-corrector method requires a derivative with respect to a scaling parameter in the interpolation, and variable projection involves differentiation through an inexact iterative least squares solve. Much of the theory for these problems exists in the literature, but perhaps due to the difficulty of some required derivatives, they do not exist in standard image registration packages like FAIR [15]. The code and data required generate the figures used in this text and to reproduce results can be found at our GitHub repository [22].
2 Image Registration Background
@@ 2.1 Numerical Optimization Framework
Assume our images are supported on a box-shaped subset $\Omega\subset\mathbb{R}^{d}$ for $d=2$ or 3.3.3 . We consider an enumerated set ${\boldsymbol{\omega}}=\{\vec{x}_{1},\dots,\vec{x}_{n}\}\subset\Omega$ of points with even spacing $h_{x}$ and $h_{y}$ in each dimension, on which we have evaluations of the reference and template images. (We do not have evaluations for points not on the grid, which we will have to address.) Using this data, we would like to be able to compute the distance $\mathcal{D}[{\cal T}\circ\vec{y},{\cal R}]$ of our choice, which in this work will be the least-squares distance metric
@@
The final thing to discretize is the transformation $\vec{y}.$ The most expressive “non-parametric” approach would be to solve for every value in $\vec{y}({\boldsymbol{\omega}})$ separately. However, to avoid such a high dimensional optimization problem, we use a parameterization in many fewer parameters ${\bf w}\in\mathbb{R}^{p}$, henceforth writing $\vec{y}(\vec{\cdot};{\bf w})$. A number of standard parameterizations are covered in [15]. In this work, we consider only affine transformations for simplicity. Our goal is to evaluate and demonstrate the behavior of the proposed methods, rather than to solve difficult image registration problems.
@@
(2) $J({\bf w})=\|{\cal T}(\vec{y}({\boldsymbol{\omega}};{\bf w}))-{\cal R}({\boldsymbol{\omega}})\|_{2}^{2}h_{x}h_{y}+\lambda^{2}S({\bf w}),$
@@ 2.2 Automatic Differentiation
To demonstrate the practical efficiency of automatic differentiation (AD), Table 1 reports average evaluation times for computing various derivatives under two transformation models: an affine transformation with 6 parameters, and a neural ODE-based transformation with 50 parameters [1, 20, 23].
We define the forward function as the map ${\bf w}\mapsto{\cal T}(\vec{y}({\boldsymbol{\omega}};{\bf w}))$, which evaluates the template image at coordinates determined by ${\bf w}$, prior to loss computation. The full loss function $J({\bf w})$ follows (2), incorporating squared differences and quadrature. All derivatives are computed using Functorch [12], which provides composable AD transforms similar to those in JAX [19].
While the theoretical cost of computing gradients and Hessians scales as $\mathcal{O}(n)$ and $\mathcal{O}(n^{2})$, respectively, modern AD frameworks such as Functorch exploit graph-level optimizations, vectorization, and batched computation to reduce overhead [12]. We observe that while the gradient and Jacobian costs scale moderately, the Hessian cost increases substantially for the 50-parameter case. This supports the practical distinction in cost between Newton and quasi-Newton methods for higher-dimensional transformations.
Table 1: Empirical cost of computing derivatives using AD. “Factor” is the ratio of runtime relative to the corresponding loss or forward function evaluation. For example, computing the gradient took 2.5× the time of a loss evaluation in the affine case. Each measurement was averaged over 100 runs. Reported values are mean runtime ($\mu$) ± standard deviation ($\sigma$) in milliseconds.
@@
Loss Function 2.171±0.2592.171$\pm$ 0.2592.171 ± 0.259 1×1$\times$1 × 5.012±2.1685.012$\pm$ 2.1685.012 ± 2.168 1×1$\times$1 ×
Gradient 5.422±0.4095.422$\pm$ 0.4095.422 ± 0.409 2.5×2.5$\times$2.5 × 13.65±1.01813.65$\pm$ 1.01813.65 ± 1.018 2.7×2.7$\times$2.7 ×
Hessian 12.44±1.79812.44$\pm$ 1.79812.44 ± 1.798 5.7×5.7$\times$5.7 × 120.0±8.737120.0$\pm$ 8.737120.0 ± 8.737 24×24$\times$24 ×
Forward Function 1.054±0.0761.054$\pm$ 0.0761.054 ± 0.076 1×1$\times$1 × 3.204±0.1493.204$\pm$ 0.1493.204 ± 0.149 1×1$\times$1 ×
Jacobian 11.59±1.51811.59$\pm$ 1.51811.59 ± 1.518 $\times1$ 54.31±3.24054.31$\pm$ 3.24054.31 ± 3.240 17×17$\times$17 ×
@@ 3.1 Derivation of Multi-scale ODE
Each step of this ODE solve requires an inverse Hessian-vector product, so it is as feasible to implement as Newton’s method for optimization.
@@ 3.2 Predictor-Corrector Method
The approach involves taking larger steps in $\theta$ when solving (4) (i.e., using coarser discretizations for the ODE). The error introduced by these larger steps is then corrected through local minimization. In practice, Newton’s method for the minimization provides rapid local convergence [17], and allows us the ability to re-use computed Hessians when taking a step in the ODE solve.
@@
(5) ${\bf w}^{*}_{\rm pred}={\bf w}^{*}(\theta_{n})+(\theta_{n}-\theta_{n+1})\left(\frac{d}{d\theta}{\bf w}^{*}(\theta_{n})\right)$
of forward Euler’s method is taken to predict the location of ${\bf w}^{*}(\theta_{n+1}),$ and is then corrected by a local minimization procedure to obtain the true ${\bf w}^{*}(\theta_{n+1}).$
Even with the predictor-corrector method, negative curvature in the local minimization problem still poses a problem. The forward Euler predictor step (5) explicitly uses the Hessian inverse, which can lead to computational instability if the Hessian is indefinite. Pure Newton steps also struggle in this scenario, since directly using indefinite Hessians can produce non-descent directions. To address this, we opt for using SciPy’s implementation of an exact trust-region Newton solver [4] for the local minimization problem, which avoids explicit Hessian inversion by solving a trust-region subproblem through eigen-decomposition [21]. A line-search Newton method is another viable alternative [17]. After convergence of the local minimization, we reuse the Hessian computed at the accepted iterate in the subsequent forward Euler predictor step, avoiding additional Hessian evaluations.
@@ 3.3 Results
Two results are shown in this section. First, the predictor-corrector method is shown to work for a non-regularized problem where registering from a single scale does not. This motivates the method as having a similar use case as the traditional multi-scale method. Then, for the predictor-corrector method, the prediction is shown to fail when using approximated Hessians via Gauss Newton, but succeeds when using Hessians computed using AD, which are exact with respect to the discretized objective (2). An implementation of Levenberg-Marquardt is used for Gauss Newton, while SciPy’s Exact Trust Region [21] is used for Newton.
@@ 3.3.1 Single Scale vs Predictor-Corrector
Consider an affine, non-regularized image registration problem with template and reference shown in the first row of Figure 2. As shown in the figure, some image registration problems benefit from using the predictor-corrector method, as opposed to optimizing from a single scale, in order to avoid getting stuck in a local minimum.
@@ 3.3.2 Approximated Hessian vs Exact Hessian
To evaluate the effectiveness of the prediction at each scale, i.e., the step of forward Euler (5) , we compare the loss and gradient at the predicted point to the loss and gradient without the prediction. This comparison effectively compares the predictor-corrector method with the traditional multi-scale method. We define the relative loss difference as $(J({\bf w}^{*})-J({\bf w}_{\rm pred}^{*}))/J({\bf w}^{*})$, and the relative gradient norm difference as $(\|\nabla J({\bf w}^{*})\|-\|\nabla J({\bf w}_{\rm pred}^{*})\|)/$ $\|\nabla J({\bf w}^{*})\|$ to give measures of how helpful the predictions are.
This experiment is illustrated in Figure 3. A positive value indicates a successful reduction in loss or gradient, whereas a negative value indicates an increase. The results show that using exact Hessians via Newton’s method tends to decrease the loss and gradient, while using approximated Hessians via Gauss-Newton tends to increase the loss and gradient. Therefore, using exact Hessians is necessary for the predictor-corrector method to provide any benefits over a traditional multi-scale method.
Figure 3: Relative loss and gradient norm, both plotted against $\theta$ on a logarithmic scale. Exact Hessians are used for blue, approximated Hessians are used for orange. All relative gradient norm difference values using approximated Hessians are less than 31000-31000- 31000.
@@ 4 Coupled Methods for Super-Resolution
We follow the framework developed in [3]. Interpolation of the reference is linear in the reference intensity vector ${\bf f}^{(0)}$, thus, we can mathematically model our transformations $\vec{y}^{(j)}$ using matrices $\mathsf{I}^{(j)}({\bf w}).$ The notation $\mathsf{I}$ with no argument is used for the identity matrix, which is consistent with viewing the identity as an application of a trivial transformation. Note that the interpolation matrix is non-linear in its dependence on the transformation parameter vector ${\bf w}.$ Our high resolution transformed templates ${\bf f}^{(j)}\in\mathbb{R}^{n},j=1,\dots,q$ are thus given by the transformations $\mathsf{I}^{(j)}({\bf w}){\bf f}^{(0)}$ of an initial reference image ${\bf f}^{(0)}$. Our observed template images ${\bf d}^{(j)}\in\mathbb{R}^{m},j=0,1\dots,q$ are modeled as ${\bf d}^{(j)}=\mathsf{K}{\bf f}^{(j)},$ where $\mathsf{K}\in\mathbb{R}^{m\times n}$ is a linear operator which lowers the image resolution ($m<n$), in our case by pooling, that is, averaging the pixels in $k\times k$ regions to reduce resolution by $k$ times in each dimension. Note that our modeling assumes that the template ${\bf d}^{(0)}$ is registered with the unknown reference ${\bf f}^{(0)}$, as the data can only ever constrain ${\bf f}^{(0)}$ up to a transformation; registering ${\bf d}^{(0)}$ with ${\bf f}^{(0)}$ eliminates this non-uniqueness and saves some computational overhead. We further let $h^{{\bf f}}_{x},h^{{\bf f}}_{y},h^{{\bf d}}_{x}$ and $h^{{\bf d}}_{y}$ be the corresponding grid sizes for the template and reference grids.
In the problem of super-resolution, we aim to reconstruct a high-resolution reference image, $\mathbf{f}^{(0)}$ from a vector of low resolution template images $\mathbf{d}=[{\mathbf{d}^{(0)}}^{\top},\dots,{\mathbf{d}^{(q)}}^{\top}]^{\top}$, which are assumed to be unregistered, down-sampled versions of $\mathbf{f}^{(0)}$. Our aim in super-resolution is to register multiple templates $\mathbf{d}^{(j)}$ to an unknown reference $\mathbf{f}^{(0)},$ while simultaneously reconstructing $\mathbf{f}^{(0)}$ from the unregistered templates.
4.1 Super-Resolution Framework
@@
$J({\bf w},{\bf f}^{(0)})=\|(\mathsf{I}\otimes\mathsf{K})\mathsf{I}({\bf w}){\bf f}^{(0)}-{\bf d}\|^{2}h^{{\bf d}}_{x}h^{{\bf d}}_{y}+\lambda_{{\bf f}}^{2}S_{{\bf f}^{(0)}}({\bf f}^{(0)}).$
@@
It should be noted that this quadratic regularizer is not ideal, as it penalizes high wavenumber content in the reconstruction that may be part of the true reference image. This shows up in our examples as “checkerboard” patterns in the residual. To move beyond this, nonlinear regularization methods such as total variation can be used with a “Linearize-And-Project” approach [11], however, like in [3], we stick to a quadratic regularization term as it simplifies both the theory and practice of variable projection.
@@ 4.2 Variable Projection Theory
We begin by noticing that (6) is in the separable form addressed by the variable projection method [9, 8, 6]. To make this even more clear, we write it in the stacked form
$J({\bf w},{\bf f}^{(0)})=\left\|\begin{bmatrix}\sqrt{h^{{\bf d}}_{x}h^{{\bf d}}_{y}}\ (\mathsf{I}\otimes\mathsf{K})\mathsf{I}({\bf w})\\ \lambda_{{\bf f}}\sqrt{h^{{\bf f}}_{x}h^{{\bf f}}_{y}}\ \mathsf{L}\end{bmatrix}{\bf f}-\begin{bmatrix}\sqrt{h^{{\bf d}}_{x}h^{{\bf d}}_{y}}\ {\bf d}\\ \mathsf{O}\end{bmatrix}\right\|^{2}=\|\mathsf{A}_{\lambda_{{\bf f}}}({\bf w}){\bf f}^{(0)}-{\bf b}\|^{2}.$
@@
(7) $\tilde{J}({\bf w})=\|\mathsf{A}_{\lambda_{{\bf f}}}({\bf w}){\bf f}^{(0)}({\bf w})-{\bf b}\|^{2}$
@@
Computing the Jacobian $\frac{\partial{\bf f}^{(0)}({\bf w})}{\partial{\bf w}}$ of the projection ${\bf f}^{(0)}({\bf w})$ requires differentiation through the least squares solution. An exact formula for such a derivative based on implicit differentiation through an exact solution exists [9], but it is expensive to compute. Previous work in variable-projected Gauss-Newton has employed approximations to avoid this cost—either through low-rank SVD approximations [16] or by entirely neglecting the Jacobian term (effectively setting $\frac{\partial{\bf f}^{(0)}({\bf w})}{\partial{\bf w}}=0$) as done in [3].
In an automatic differentiation framework, we can ignore the projection term by turning gradient tracking off during the least squares solve. However, by choosing to track gradients through the least squares solution, we can include this term at the cost of increased memory overhead. It has been shown that automatic differentiation through the Conjugate Gradient (CG) algorithm gives correct gradients in the case of an inexact solve [10], and in the case of an exact solve [2], which we will use in our solutions. A recent preprint [13] explores automatic differentiation through other least-squares algorithms.